# The Effect of Intrinsic Dataset Properties on Generalization: Unraveling Learning Differences Between Natural and Medical Images

### 16 Jan 2024 v1

@@ 1 Introduction
First, we introduce the novel measure of the intrinsic label sharpness ($K_{\mathcal{F}}$) of a dataset (defined in Section 3.2). The label sharpness essentially measures how similar images in the dataset can be to each other while still having different labels, and we find that it usually differs noticeably between natural and medical image datasets. We then derive and test a neural network generalization scaling law with respect to dataset intrinsic dimension $d_{\mathrm{data}}$, which includes $K_{\mathcal{F}}$. Our experiments support the derived scaling behavior within each of these two domains, and show a distinct difference in the scaling rate between them. According to our scaling law, this may be due to the measured $K_{\mathcal{F}}$ being typically higher for medical datasets.
Next, we show how a model’s adversarial robustness relates to its training set’s $K_{\mathcal{F}}$, and show that over a range of attacks, robustness decreases with higher $K_{\mathcal{F}}$. Medical image datasets, which have higher $K_{\mathcal{F}}$, are typically more susceptible to adversarial attack than natural image datasets. Finally, we extend our $d_{\mathrm{data}}$ formalism to derive and test a generalization scaling law with respect to the intrinsic dimension of the model’s learned representations, $d_{\mathrm{repr}}$, and reconcile the $d_{\mathrm{data}}$ and $d_{\mathrm{repr}}$ scaling laws to show that $d_{\mathrm{data}}$ serves as an approximate upper bound for $d_{\mathrm{repr}}$. We also provide many additional results in the supplementary material, such as a likelihood analysis of our proposed scaling law given observed generalization data, the evaluation of a new dataset in a third domain, an example of a practical application of our findings, and more.
All theoretical results are validated with thorough experiments on six convolutional models and eleven datasets from natural and medical imaging domains over a range of training set sizes. We hope that our work initiates further study into how network behavior differs between imaging domains.
@@ 3 Preliminaries
As in Bahri et al. (2021), we consider over-parameterized (number of parameters $\gg N$) models $f(x):\mathbb{R}^{n}\rightarrow\{0,1\}$, that are “well-trained” and learn to interpolate all training data: $f(x)=\mathcal{F}(x)$ for all $x\in\mathcal{D}_{\mathrm{train}}$. We use a non-negative loss function $L$, such that $L=0$ when $f(x)=\mathcal{F}(x)$. Note that we write $L$ as the expected loss over a set of test set points. We assume that $\mathcal{F}$, $f$ and $L$ are Lipschitz/smooth on $\mathcal{M}_{d_{\mathrm{data}}}$ with respective constants $K_{\mathcal{F}}$, $K_{f}$ and $K_{L}$. Note that we use the term “Lipschitz constant” of a function to refer to the smallest value that satisfies the Lipschitz inequality. [^3] A subtlety here is that our Lipschitz assumptions only involve pairs of neighboring datapoints sampled from the true data manifold $\mathcal{M}_{d_{\mathrm{data}}}$; adversarially-perturbed images (Goodfellow et al., 2015) are not included. We focus on binary classification as in Pope et al. (2020); Konz et al. (2022), but we note that our results extend naturally to the multi-class case (see Appendix A.1 for more details).
@@ 5.2 Generalization Discrepancies Between Imaging Domains
We see that within an imaging domain (natural or medical), model generalization error typically increases with $d_{\mathrm{data}}$, as predicted, similar to prior results (Pope et al., 2020; Konz et al., 2022); in particular, approximately $\log L\propto-1/d_{\mathrm{data}}+\mathrm{const.}$, aligning with Eq. (2). However, we also see that the generalization error scaling is much sharper for models trained on medical data than natural data; models trained on datasets with similar $d_{\mathrm{data}}$ and of the same size $N$ tend to perform much worse if the data is medical images. A similarly large gap appears for the scaling of test accuracy (Appendix E.1). We posit that this difference is explained by medical datasets typically having much higher label sharpness ($\hat{K}_{\mathcal{F}}\sim 2.5\times 10^{-4}$) than natural images ($\hat{K}_{\mathcal{F}}\sim 1\times 10^{-4}$) (Fig. 1) , as $K_{\mathcal{F}}$ is the only term in Eq. (2) that differs between two models with the same training set intrinsic dimension $d_{\mathrm{data}}$ and size $N$. Moreover, in Appendix C.1 we show that accounting for $K_{\mathcal{F}}$ increases the likelihood of the posited scaling law given the observed data. However, we note that there could certainly be other factors causing the discrepancy which are not accounted for.
@@ 7 Connecting Representation Intrinsic Dimension to Dataset Intrinsic Dimension and Generalization
Our findings provide insights into how neural network behavior varies within and between the two crucial domains of natural and medical images, enhancing our understanding of generalization, representation learning, and robustness.

### 21 Feb 2024 v3

@@ 1 Introduction
First, we introduce the novel measure of the intrinsic label sharpness ($K_{\mathcal{F}}$) of a dataset (defined in Section 3.2). The label sharpness essentially measures how similar images in the dataset can be to each other while still having different labels, and we find that it usually differs noticeably between natural and medical image datasets. We then derive and test a neural network generalization scaling law with respect to dataset intrinsic dimension $d_{\mathrm{data}}$, which includes $K_{\mathcal{F}}$. Our experiments support the derived scaling behavior within each of these two domains, and show a distinct difference in the scaling rate between them. According to our scaling law and likelihood analysis of observed generalization data (Appendix C.1), this may be due to the measured $K_{\mathcal{F}}$ being typically higher for medical datasets.
Next, we show how a model’s adversarial robustness relates to its training set’s $K_{\mathcal{F}}$, and show that over a range of attacks, robustness decreases with higher $K_{\mathcal{F}}$. Indeed, medical image datasets, which have higher $K_{\mathcal{F}}$, are typically more susceptible to adversarial attack than natural image datasets. Finally, we extend our $d_{\mathrm{data}}$ formalism to derive and test a generalization scaling law with respect to the intrinsic dimension of the model’s learned representations, $d_{\mathrm{repr}}$, and reconcile the $d_{\mathrm{data}}$ and $d_{\mathrm{repr}}$ scaling laws to show that $d_{\mathrm{data}}$ serves as an approximate upper bound for $d_{\mathrm{repr}}$. We also provide many additional results in the supplementary material, such as a likelihood analysis of our proposed scaling law given observed generalization data (Appendix C.1), the evaluation of a new dataset in a third domain (Appendix C.2), an example of a practical application of our findings (Appendix C.3), and more.
All theoretical results are validated with thorough experiments on six CNN architectures and eleven datasets from natural and medical imaging domains over a range of training set sizes. We hope that our work initiates further study into how network behavior differs between imaging domains.
@@ 3 Preliminaries
As in Bahri et al. (2021), we consider over-parameterized (number of parameters $\gg N$) models $f(x):\mathbb{R}^{n}\rightarrow\{0,1\}$, that are “well-trained” and learn to interpolate all training data: $f(x)=\mathcal{F}(x)$ for all $x\in\mathcal{D}_{\mathrm{train}}$. We use a non-negative loss function $L$, such that $L=0$ when $f(x)=\mathcal{F}(x)$. Note that we write $L$ as the expected loss over a set of test set points. We assume that $\mathcal{F}$, $f$ and $L$ are Lipschitz/smooth on $\mathcal{M}_{d_{\mathrm{data}}}$ with respective constants $K_{\mathcal{F}}$, $K_{f}$ and $K_{L}$. Note that we use the term “Lipschitz constant” of a function to refer to the smallest value that satisfies the Lipschitz inequality. [^3] A subtlety here is that our Lipschitz assumptions only involve pairs of datapoints sampled from the true data manifold $\mathcal{M}_{d_{\mathrm{data}}}$; adversarially-perturbed images (Goodfellow et al., 2015) are not included. We focus on binary classification as in Pope et al. (2020); Konz et al. (2022), but we note that our results extend naturally to the multi-class case (see Appendix A.1 for more details).
@@ 5.2 Generalization Discrepancies Between Imaging Domains
We see that within an imaging domain (natural or medical), model generalization error typically increases with $d_{\mathrm{data}}$, as predicted, similar to prior results (Pope et al., 2020; Konz et al., 2022); in particular, approximately $\log L\propto-1/d_{\mathrm{data}}+\mathrm{const.}$, aligning with Eq. (2). However, we also see that the generalization error scaling is much sharper for models trained on medical data than natural data; models trained on datasets with similar $d_{\mathrm{data}}$ and of the same size $N$ tend to perform much worse if the data is medical images. A similarly large gap appears for the scaling of test accuracy (Appendix E.1). We posit that this difference is explained by medical datasets typically having much higher label sharpness ($\hat{K}_{\mathcal{F}}\sim 2.5\times 10^{-4}$) than natural images ($\hat{K}_{\mathcal{F}}\sim 1\times 10^{-4}$) (Fig. 1) , as $K_{\mathcal{F}}$ is the only term in Eq. (2) that differs between two models with the same training set intrinsic dimension $d_{\mathrm{data}}$ and size $N$. Moreover, in Appendix C.1 we show that accounting for $K_{\mathcal{F}}$ increases the likelihood of the posited scaling law given the observed generalization data. However, we note that there could certainly be other factors causing the discrepancy which are not accounted for.
@@ 7 Connecting Representation Intrinsic Dimension to Dataset Intrinsic Dimension and Generalization
Our findings provide insights into how neural network behavior varies within and between the two crucial domains of natural and medical images, enhancing our understanding of the dependence of generalization ability, representation learning, and adversarial robustness on intrinsic measurable properties of the training set.
@@ A.1 Extension of Results to Multi-Class Classification
Label sharpness.
The label sharpness metric $\hat{K}_{\mathcal{F}}$ (Eq. 1) was formulated under the binary classification scenario, where data is either labeled with 00 or $1$ (Sec. 3). However, it could potentially be extended to the multi-class scenario by simply replacing the $|y_{j}-y_{k}|$ term in the numerator of Eq. 1 with the indicator function $1_{y_{j}\neq y_{k}}$ as
$\hat{K}_{\mathcal{F}}:=\max\limits_{j,k}\left(\frac{1_{y_{j}\neq y_{k}}}{||x_{j}-x_{k}||}\right),$ (6)
which clearly simplifies to Eq. 1 for binary classification. This modification prevents $\hat{K}_{\mathcal{F}}$ from being biased by the numerical value of labels given to different classes, but a more careful extension could be pursued in the future to confirm a properly theoretically-motivated multi-class label sharpness metric.